# Danny Halawi4, Fred Zhang4, Philip E. Tetlock2,5

### v2

@@ 6 Discussion
[…]
From a practical point of view, we hope that our benchmark can shed light on how automated systems compare in accuracy to humans and when they might become the preferred source of predictions. Finally, from the perspective of LLM evaluation, forecasting requires models to integrate pre-training knowledge, engage in effective reasoning, and maintain calibration to avoid overconfidence. All of these are essential capabilities that are desired across various domains. We believe that performance on our benchmark can shed light on other capabilities of LLMs.
@@ B.1 Question and resolution sets
The date the forecast was due in ISO 8601 format YYYY-MM-DD; day $D$ from Figure 3. string
@@ B.2.1 General public forecast set
reasoning The rationale underlying the forecast. string
@@ B.2.2 Superforecaster forecast set
In addition to the fields outlined in Table 10 and Table 11, the superforecaster dataset contains forecasts with the fields shown in Table 12. Likewise, this dataset will be updated every time a group of superforecasters is surveyed.
@@ D.1 Instructions
Participants in the public survey were first prompted to read a consent form detailing the tasks involved in the study and potential risks to participants, alongside estimates of payment for completing each study (\$5 + bonus payments for forecasting performance in the introductory survey; \$10 for completing the primary survey). Then, participants were prompted to give their responses to 20 forecasting questions randomly selected from the 200-question subset of questions provided to the LLMs.
@@ D.2 Screenshots
Figure 5: An example question generated from a data provider, in this case DBnomics, from the public survey. Two of eight time horizons for which we elicited forecasts are included above.
@@ Appendix F Prompts
In this section, we present the following prompts: zero-shot (Figure 6), scratchpad (Figure 7), and three prompts (Figure 8, Figure 9, Figure 10) written by superforecasters from the Forecasting Research Institute used to construct the LLM "crowd" baseline. Note that the scratchpad with information retrieval prompt is simply the scratchpad with an additional line "We have retrieved the following information for this question: {retrieved_info}" before the instructions begin. For combination questions, we slightly modify each of the above prompts by including the second question’s information (Figure 11).
@@ Appendix G Statistical Details
Appendix H Top 50 ForecastBench Leaderboards
We show the Top 50 leaderboard with human performance evaluated on 200 questions received by humans in Table 17 and the one excluding humans evaluated on the full 1,000 questions in Table 18. Additionally, we show the leaderboard with the inclusion of combination questions in Table 19 , where we generate human forecasts on the combination questions by treating the human forecasts on each component of a combination question as independent (multipling them together to produce the forecast on the relevant combination question). Lastly, we show the leaderboard with combination questions Table 20, where we used the same human-generated forecasts as above but also generated combination forecasts for LLMs from their forecasts on the individual questions, assuming independence of their forecasts on each component, and ignoring the forecast they actually gave on the combination question.
Table 17: Leaderboard with Human Performance: Top 50
% Brier Score $\downarrow$ Model Organization Information provided Prompt Dataset (N=210) Market (N=77) Overall (N=287) Confidence Interval Pairwise p-value comparing to No. 1 Pct. more accurate than No. 1 Superforecaster median forecast ForecastBench 0.135 0.051 0.093 [0.073, 0.112] 0% Public median forecast ForecastBench 0.167 0.048 0.107 [0.09, 0.125] 0.009 30% Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.155 0.067 0.111 [0.09, 0.132] 0.01 33% GPT-4 OpenAI Freeze values Zero shot 0.169 0.056 0.113 [0.093, 0.132] 0.003 37% Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.155 0.074 0.114 [0.09, 0.139] 0.008 36% Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.16 0.07 0.115 [0.095, 0.135] 0.004 31% GPT-4o OpenAI Freeze values Scratchpad 0.173 0.059 0.116 [0.098, 0.133] 0.002 30% GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.177 0.055 0.116 [0.097, 0.135] <0.001 37% GPT-4o OpenAI News with freeze values Scratchpad 0.179 0.059 0.119 [0.101, 0.137] <0.001 28% Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.163 0.081 0.122 [0.107, 0.137] <0.001 24% Mistral-Large-Latest Mistral AI Freeze values Zero shot 0.173 0.074 0.123 [0.102, 0.145] <0.001 28% Claude-3-5-Sonnet-20240620 Anthropic No Scratchpad 0.155 0.093 0.124 [0.103, 0.145] <0.001 29% Llama-3-70b-Chat-Hf Meta Freeze values Zero shot 0.175 0.076 0.125 [0.104, 0.146] <0.001 30% Claude-3-5-Sonnet-20240620 Anthropic News No 0.171 0.082 0.126 [0.108, 0.145] <0.001 28% Claude-3-5-Sonnet-20240620 Anthropic News No 0.167 0.089 0.128 [0.107, 0.148] <0.001 28% Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.185 0.071 0.128 [0.106, 0.151] <0.001 33% GPT-4 OpenAI Freeze values Scratchpad 0.188 0.069 0.128 [0.114, 0.143] <0.001 24% Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.16 0.097 0.128 [0.108, 0.149] <0.001 26% GPT-4o OpenAI No Scratchpad 0.173 0.084 0.128 [0.111, 0.146] <0.001 26% Claude-2.1 Anthropic Freeze values Scratchpad 0.206 0.055 0.131 [0.116, 0.145] <0.001 30% Qwen1.5-110B-Chat Qwen Freeze values Zero shot 0.196 0.065 0.131 [0.113, 0.149] <0.001 29% Llama-3-70b-Chat-Hf Meta Freeze values Scratchpad 0.188 0.076 0.132 [0.117, 0.147] <0.001 24% Gemini-1.5-Pro Google Freeze values Zero shot 0.19 0.076 0.133 [0.109, 0.157] <0.001 31% GPT-4o OpenAI Freeze values Zero shot 0.205 0.063 0.134 [0.112, 0.156] <0.001 35% Claude-3-Opus-20240229 Anthropic Freeze values Scratchpad 0.176 0.092 0.134 [0.115, 0.152] <0.001 26% GPT-4o OpenAI News Scratchpad 0.179 0.089 0.134 [0.113, 0.155] <0.001 25% Claude-3-5-Sonnet-20240620 Anthropic No Zero shot 0.155 0.114 0.134 [0.109, 0.16] <0.001 28% Gemini-1.5-Flash Google Freeze values Scratchpad 0.195 0.077 0.136 [0.117, 0.155] <0.001 25% Gemini-1.5-Flash Google Freeze values Zero shot 0.183 0.089 0.136 [0.11, 0.163] <0.001 30% GPT-4-Turbo-2024-04-09 OpenAI No Zero shot 0.177 0.098 0.137 [0.118, 0.156] <0.001 22% Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Zero shot 0.195 0.079 0.137 [0.114, 0.16] <0.001 32% Claude-3-Opus-20240229 Anthropic No Scratchpad 0.176 0.1 0.138 [0.121, 0.155] <0.001 24% GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.191 0.085 0.138 [0.116, 0.16] <0.001 27% Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Scratchpad 0.19 0.086 0.138 [0.122, 0.154] <0.001 25% Claude-2.1 Anthropic No Scratchpad 0.206 0.07 0.138 [0.121, 0.156] <0.001 28% Gemini-1.5-Pro Google News with freeze values Scratchpad 0.181 0.096 0.138 [0.12, 0.157] <0.001 25% Gemini-1.5-Pro Google Freeze values Scratchpad 0.179 0.099 0.139 [0.121, 0.157] <0.001 26% GPT-4 OpenAI No Scratchpad 0.188 0.096 0.142 [0.128, 0.157] <0.001 20% Mistral-Large-Latest Mistral AI No Scratchpad 0.163 0.123 0.143 [0.126, 0.16] <0.001 24% GPT-4-Turbo-2024-04-09 OpenAI No Scratchpad 0.191 0.095 0.143 [0.126, 0.16] <0.001 23% Imputed Forecaster ForecastBench 0.25 0.037 0.143 [0.129, 0.158] <0.001 34% Claude-2.1 Anthropic Freeze values Zero shot 0.22 0.067 0.144 [0.124, 0.163] <0.001 31% Gemini-1.5-Pro Google No Scratchpad 0.179 0.109 0.144 [0.125, 0.163] <0.001 24% GPT-4 OpenAI No Zero shot 0.169 0.119 0.144 [0.127, 0.162] <0.001 23% Gemini-1.5-Pro Google News Scratchpad 0.181 0.112 0.146 [0.127, 0.166] <0.001 26% GPT-4-Turbo-2024-04-09 OpenAI News with freeze values Scratchpad 0.202 0.09 0.146 [0.124, 0.169] <0.001 31% Llama-3-70b-Chat-Hf Meta No Zero shot 0.175 0.12 0.148 [0.128, 0.168] <0.001 24% Claude-3-Opus-20240229 Anthropic No Zero shot 0.185 0.112 0.149 [0.125, 0.173] <0.001 25% Qwen1.5-110B-Chat Qwen News with freeze values Scratchpad 0.196 0.103 0.149 [0.132, 0.167] <0.001 25% Mixtral-8x7B-Instruct-V0.1 Mistral AI Freeze values Zero shot 0.209 0.09 0.15 [0.12, 0.179] <0.001 36%
Notes:
% reduces space between this line and the next one Pairwise p-value comparing to No. 1 (bootstrapped): The p-value calculated by bootstrapping the differences in overall score between each model and the best forecaster (the group with rank 1) under the null hypothesis that there’s no difference.
% reduces space between this line and the next one Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster (with rank 1).
Table 18: Leaderboard excluding Human Performance: Top 50
% Brier Score $\downarrow$ Model Organization Information provided Prompt Dataset (N=210) Market (N=77) Overall (N=287) Confidence Interval Pairwise p-value comparing to No. 1 Pct. more accurate than No. 1 GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.175 0.051 0.113 [0.107, 0.118] 0% Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.179 0.047 0.113 [0.108, 0.119] 0.442 53% GPT-4o OpenAI Freeze values Scratchpad 0.201 0.039 0.12 [0.115, 0.125] <0.001 45% Gemini-1.5-Pro Google Freeze values Scratchpad 0.168 0.075 0.121 [0.116, 0.127] <0.001 42% GPT-4o OpenAI News with freeze values Scratchpad 0.196 0.05 0.123 [0.117, 0.128] <0.001 41% Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.192 0.061 0.127 [0.12, 0.133] <0.001 44% Gemini-1.5-Pro Google News with freeze values Scratchpad 0.172 0.082 0.127 [0.121, 0.132] <0.001 41% GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.199 0.055 0.127 [0.12, 0.133] <0.001 47% Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.176 0.078 0.127 [0.122, 0.133] <0.001 32% GPT-4-Turbo-2024-04-09 OpenAI No Scratchpad 0.175 0.083 0.129 [0.124, 0.134] <0.001 7% Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.197 0.067 0.132 [0.127, 0.138] <0.001 43% Gemini-1.5-Pro Google No Scratchpad 0.168 0.097 0.133 [0.127, 0.138] <0.001 39% GPT-4 OpenAI Freeze values Scratchpad 0.198 0.067 0.133 [0.127, 0.138] <0.001 33% Claude-3-5-Sonnet-20240620 Anthropic No Scratchpad 0.179 0.087 0.133 [0.127, 0.139] <0.001 46% Llama-3-70b-Chat-Hf Meta Freeze values Zero shot 0.199 0.069 0.134 [0.128, 0.14] <0.001 38% Gemini-1.5-Pro Google News Scratchpad 0.172 0.097 0.134 [0.128, 0.14] <0.001 39% Gemini-1.5-Pro Google Freeze values Zero shot 0.219 0.052 0.135 [0.129, 0.141] <0.001 44% Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.208 0.063 0.136 [0.129, 0.142] <0.001 51% Imputed Forecaster ForecastBench 0.25 0.021 0.136 [0.133, 0.138] <0.001 43% LLM Crowd ForecastBench News 0.234 0.039 0.137 [0.134, 0.14] <0.001 37% GPT-4-Turbo-2024-04-09 OpenAI News with freeze values Scratchpad 0.213 0.061 0.137 [0.132, 0.142] <0.001 29% Gemini-1.5-Flash Google Freeze values Scratchpad 0.187 0.088 0.138 [0.131, 0.144] <0.001 40% GPT-4o OpenAI News Scratchpad 0.196 0.082 0.139 [0.133, 0.145] <0.001 34% Qwen1.5-110B-Chat Qwen No Scratchpad 0.176 0.102 0.139 [0.134, 0.144] <0.001 30% LLM Crowd ForecastBench News 0.236 0.042 0.139 [0.136, 0.142] <0.001 36% LLM Crowd ForecastBench News 0.237 0.042 0.139 [0.136, 0.142] <0.001 36% GPT-4 OpenAI Freeze values Zero shot 0.222 0.057 0.14 [0.134, 0.145] <0.001 38% Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.197 0.082 0.14 [0.135, 0.145] <0.001 29% GPT-4o OpenAI No Scratchpad 0.201 0.081 0.141 [0.136, 0.147] <0.001 37% Llama-3-70b-Chat-Hf Meta Freeze values Scratchpad 0.208 0.075 0.142 [0.137, 0.147] <0.001 29% GPT-4-Turbo-2024-04-09 OpenAI No Zero shot 0.199 0.086 0.142 [0.136, 0.148] <0.001 38% Gemini-1.5-Pro Google No Zero shot 0.219 0.069 0.144 [0.138, 0.15] <0.001 42% Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.197 0.091 0.144 [0.139, 0.15] <0.001 40% GPT-4 OpenAI No Scratchpad 0.198 0.091 0.145 [0.14, 0.149] <0.001 29% Claude-3-Opus-20240229 Anthropic Freeze values Scratchpad 0.21 0.08 0.145 [0.139, 0.151] <0.001 36% Mistral-Large-Latest Mistral AI Freeze values Zero shot 0.201 0.091 0.146 [0.139, 0.153] <0.001 35% Gemini-1.5-Pro Google News No 0.193 0.099 0.146 [0.14, 0.152] <0.001 34% Gemini-1.5-Flash Google No Scratchpad 0.187 0.106 0.147 [0.141, 0.153] <0.001 37% Gemini-1.5-Flash Google Freeze values Zero shot 0.225 0.068 0.147 [0.14, 0.154] <0.001 45% Llama-3-70b-Chat-Hf Meta No Zero shot 0.199 0.095 0.147 [0.141, 0.152] <0.001 34% Claude-3-5-Sonnet-20240620 Anthropic News No 0.199 0.096 0.148 [0.142, 0.153] <0.001 38% Claude-3-Opus-20240229 Anthropic No Zero shot 0.192 0.104 0.148 [0.141, 0.155] <0.001 37% Qwen1.5-110B-Chat Qwen News with freeze values Scratchpad 0.207 0.09 0.149 [0.143, 0.154] <0.001 27% GPT-4-Turbo-2024-04-09 OpenAI News Scratchpad 0.213 0.085 0.149 [0.144, 0.154] <0.001 24% GPT-4-Turbo-2024-04-09 OpenAI News No 0.212 0.086 0.149 [0.144, 0.154] <0.001 27% Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Scratchpad 0.214 0.091 0.152 [0.146, 0.158] <0.001 28% Claude-2.1 Anthropic No Scratchpad 0.232 0.072 0.152 [0.148, 0.157] <0.001 36% GPT-4o OpenAI News No 0.212 0.093 0.152 [0.147, 0.158] <0.001 28% Claude-3-5-Sonnet-20240620 Anthropic No Zero shot 0.208 0.1 0.154 [0.147, 0.161] <0.001 43% Claude-3-5-Sonnet-20240620 Anthropic News No 0.213 0.096 0.154 [0.149, 0.16] <0.001 34%
Notes:
% reduces space between this line and the next one Pairwise p-value comparing to No. 1 (bootstrapped): The p-value calculated by bootstrapping the differences in overall score between each model and the best forecaster (the group with rank 1) under the null hypothesis that there’s no difference.
% reduces space between this line and the next one Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster (with rank 1).
Table 19: Leaderboard with Performance om Combination Questions: Top 50
% Brier Score $\downarrow$ Model Organization Information provided Prompt Dataset (N=210) Market (N=77) Overall (N=287) Confidence Interval Pairwise p-value comparing to No. 1 Pct. more accurate than No. 1 Superforecaster median forecast ForecastBench 0.097 0.04 0.068 [0.06, 0.077] 0% Public median forecast ForecastBench 0.127 0.039 0.083 [0.075, 0.091] <0.001 28% Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.168 0.069 0.119 [0.107, 0.13] <0.001 27% GPT-4o OpenAI Freeze values Scratchpad 0.191 0.048 0.119 [0.11, 0.129] <0.001 28% GPT-4o OpenAI News with freeze values Scratchpad 0.179 0.06 0.119 [0.11, 0.129] <0.001 27% GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.171 0.068 0.12 [0.109, 0.131] <0.001 27% Gemini-1.5-Pro Google Freeze values Scratchpad 0.16 0.092 0.126 [0.116, 0.137] <0.001 24% GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.193 0.064 0.128 [0.117, 0.139] <0.001 30% Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.173 0.088 0.131 [0.119, 0.142] <0.001 23% Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.183 0.083 0.133 [0.12, 0.145] <0.001 27% Gemini-1.5-Pro Google News with freeze values Scratchpad 0.161 0.104 0.133 [0.122, 0.144] <0.001 22% Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.173 0.093 0.133 [0.123, 0.143] <0.001 18% GPT-4-Turbo-2024-04-09 OpenAI No Scratchpad 0.171 0.096 0.134 [0.124, 0.143] <0.001 19% Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.189 0.079 0.134 [0.122, 0.147] <0.001 26% GPT-4o OpenAI News Scratchpad 0.179 0.093 0.136 [0.125, 0.147] <0.001 21% Gemini-1.5-Pro Google No Scratchpad 0.16 0.112 0.136 [0.126, 0.147] <0.001 21% GPT-4 OpenAI Freeze values Scratchpad 0.192 0.08 0.136 [0.126, 0.146] <0.001 21% Claude-3-5-Sonnet-20240620 Anthropic No Scratchpad 0.168 0.105 0.137 [0.125, 0.148] <0.001 21% Llama-3-70b-Chat-Hf Meta Freeze values Zero shot 0.188 0.089 0.138 [0.127, 0.15] <0.001 23% GPT-4 OpenAI Freeze values Zero shot 0.213 0.064 0.139 [0.128, 0.149] <0.001 27% Gemini-1.5-Pro Google Freeze values Zero shot 0.208 0.069 0.139 [0.127, 0.15] <0.001 26% GPT-4o OpenAI No Scratchpad 0.191 0.086 0.139 [0.128, 0.149] <0.001 20% Gemini-1.5-Pro Google News Scratchpad 0.161 0.116 0.139 [0.127, 0.15] <0.001 21% Imputed Forecaster ForecastBench 0.25 0.029 0.14 [0.134, 0.145] <0.001 29% LLM Crowd ForecastBench News 0.232 0.048 0.14 [0.134, 0.146] <0.001 24% Gemini-1.5-Flash Google Freeze values Scratchpad 0.175 0.107 0.141 [0.129, 0.153] <0.001 22% Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.173 0.11 0.142 [0.13, 0.153] <0.001 20% LLM Crowd ForecastBench News 0.234 0.051 0.142 [0.136, 0.148] <0.001 24% LLM Crowd ForecastBench News 0.236 0.05 0.143 [0.137, 0.149] <0.001 24% Llama-3-70b-Chat-Hf Meta Freeze values Scratchpad 0.196 0.089 0.143 [0.134, 0.152] <0.001 18% Claude-3-Opus-20240229 Anthropic Freeze values Scratchpad 0.192 0.095 0.143 [0.133, 0.154] <0.001 19% Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.186 0.102 0.144 [0.134, 0.153] <0.001 18% Gemini-1.5-Flash Google No Scratchpad 0.175 0.114 0.144 [0.134, 0.155] <0.001 17% Claude-3-5-Sonnet-20240620 Anthropic News No 0.178 0.111 0.145 [0.133, 0.156] <0.001 19% GPT-4-Turbo-2024-04-09 OpenAI News with freeze values Scratchpad 0.21 0.08 0.145 [0.134, 0.156] <0.001 24% Qwen1.5-110B-Chat Qwen No Scratchpad 0.173 0.119 0.146 [0.137, 0.155] <0.001 16% GPT-4-Turbo-2024-04-09 OpenAI No Zero shot 0.193 0.1 0.146 [0.134, 0.158] <0.001 21% Gemini-1.5-Pro Google No Zero shot 0.208 0.087 0.148 [0.136, 0.16] <0.001 21% GPT-4 OpenAI No Scratchpad 0.192 0.107 0.149 [0.141, 0.158] <0.001 16% Claude-2.1 Anthropic No Scratchpad 0.224 0.076 0.15 [0.141, 0.159] <0.001 24% Gemini-1.5-Flash Google Freeze values Zero shot 0.216 0.084 0.15 [0.136, 0.163] <0.001 27% Mistral-Large-Latest Mistral AI Freeze values Zero shot 0.197 0.104 0.15 [0.138, 0.163] <0.001 22% Gemini-1.5-Pro Google News No 0.181 0.122 0.151 [0.14, 0.163] <0.001 20% Claude-3-Opus-20240229 Anthropic No Zero shot 0.183 0.122 0.152 [0.139, 0.165] <0.001 22% GPT-4-Turbo-2024-04-09 OpenAI News Scratchpad 0.21 0.097 0.153 [0.143, 0.164] <0.001 19% Qwen1.5-110B-Chat Qwen News with freeze values Scratchpad 0.199 0.109 0.154 [0.144, 0.165] <0.001 18% Llama-3-70b-Chat-Hf Meta No Zero shot 0.188 0.121 0.154 [0.144, 0.165] <0.001 17% Claude-2.1 Anthropic Freeze values Scratchpad 0.224 0.089 0.156 [0.146, 0.167] <0.001 23% GPT-4-Turbo-2024-04-09 OpenAI News No 0.206 0.108 0.157 [0.147, 0.168] <0.001 19% Mistral-Large-Latest Mistral AI No Scratchpad 0.186 0.131 0.158 [0.149, 0.168] <0.001 16%
Notes:
% reduces space between this line and the next one Pairwise p-value comparing to No. 1 (bootstrapped): The p-value calculated by bootstrapping the differences in overall score between each model and the best forecaster (the group with rank 1) under the null hypothesis that there’s no difference.
% reduces space between this line and the next one Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster (with rank 1).
Table 20: Leaderboard with Performance on Combination Questions (Independence): Top 50

### v3

@@ 6 Discussion
We introduced ForecastBench, a dynamic and continuously updated benchmark for evaluating automated forecasting systems. By focusing exclusively on questions that are unresolved at the time of submission, we eliminate the risks of data leakage and ensure a robust evaluation environment. Our initial results demonstrate that while state-of-the-art LLMs exhibit promising potential, they underperform expert human forecasters. This performance gap highlights the challenges in leveraging current LLMs for accurate, real-time forecasting.
We produce a public leaderboard listing the real-time accuracy of top LLMs and humans as well as a standardized dataset of forecasting questions and rationales. Future work should leverage this auxiliary dataset of predictions and rationales to fine-tune models, explore new architectures, and develop adaptive systems better suited for general reasoning in dynamic, real-world environments. Ultimately, ForecastBench serves as a step toward harnessing the full potential of AI-based systems and forecasting in decision-making.
@@ B.1 Question and resolution sets
The date the forecast is due in ISO 8601 format YYYY-MM-DD. string
@@ B.2.1 General public forecast set
reasoning The rationale underlying the forecast. During data anonymization, we insert [redacted to maintain anonymity] wherever text has been redacted. string
@@
suspected_LLM true if we suspect the rationale was created by an LLM, false otherwise. boolean
@@ B.2.2 Superforecaster forecast set
In addition to the fields outlined in Table 11 and Table 12, the superforecaster dataset contains forecasts with the fields shown in Table 13 (minus the suspected_LLM field). Likewise, this dataset will be updated every time a group of superforecasters is surveyed.
@@ D.1 Instructions
Participants in the public survey were first prompted to read a consent form detailing the tasks involved in the study and potential risks to participants, alongside estimates of payment for completing each study (\$5 + bonus payments for forecasting performance in the introductory survey; \$10 for completing the primary survey). In said consent form, the overall task was described as follows: "You will be asked to read some material, to follow some instructions, and to provide forecasts for future events." Then, participants were prompted to give their responses to 20 forecasting questions randomly selected from the 200-question subset of questions provided to the LLMs.
@@
The public survey was hosted on Qualtrics’s (https://www.qualtrics.com/) survey software platform.
@@
The superforecaster survey was hosted on Quorum (https://quorumapp.com/), a platform designed to host multi-stage forecasting tournaments.
@@ D.2 Screenshots
Figure 3: An example question generated from a data provider, in this case DBnomics, from the public survey. Two of eight forecast horizons for which we elicited forecasts are included above. The rationale text boxes (one for each forecast horizon) have also been excluded from the screenshot for brevity.
@@ Appendix F Prompts
In this section, we present the following prompts: zero-shot (Figure 4), scratchpad (Figure 5), and three prompts (Figure 6, Figure 7, Figure 8) written by superforecasters from the Forecasting Research Institute used to construct the LLM "crowd" baseline. Note that the scratchpad with information retrieval prompt is simply the scratchpad with an additional line "We have retrieved the following information for this question: {retrieved_info}" before the instructions begin. For combination questions, we slightly modify each of the above prompts by including the second question’s information (Figure 9). Finally, we present the prompt used to determine whether a question is marked as valid or invalid in Figure 10.
@@
{mdframed}
I want to assess the quality of a forecast question.Here is the forecast question: {question}.Please flag questions that don’t seem appropriate by outputting "flag". Otherwise, if it seems like a reasonable question or if you’re unsure, output "ok."In general, poorly-defined questions, questions that are sexual in nature, questions that are too personal, questions about the death/life expectancy of an individual should be flagged or, more generally, questions that are not in the public interest should be flagged. Geopolitical questions, questions about court cases, the entertainment industry, wars, public figures, and, more generally, questions in the public interest should be marked as "ok."Examples of questions that should be flagged:* "Will I finish my homework tonight?"* "Metaculus party 2023"* "Will Hell freeze over?"* "Heads or tails?"* "Will I get into MIT?"* "Will this video reach 100k views by the EOD?"* "If @Aella goes on the Whatever podcast, will she regret it?"* "Daily coinflip"* "Musk vs Zuckerberg: Will either of them shit their pants on the mat?"Examples of questions that should NOT be flagged:* "Will Megan Markle and Prince Harry have a baby by the end of the year?"* "Will the Brain Preservation Foundation’s Large Mammal preservation prize be won by Feb 9th, 2017?"* "Will there be more novel new drugs approved by the FDA in 2016 than in 2015?"* "Will Israel invade Rafah in May 2024?"* "Will Iraq return its ambassador to Iran in the next month?"* "Tiger Woods Will Win Another PGA Tournament"* "Will Dwayne Johnson win the 2024 US Presidential Election?"* "Will Oppenheimer win best picture AND Bitcoin reach \$70K AND Nintendo announce a new console by EOY 2024?"* "Will anybody born before 2000 live to be 150?"* "Will Taylor Swift get married before Bitcoin reaches \$100K USD?"* "Will Russia’s total territory decrease by at least 20% before 2028?"* "Will Donald Trump be jailed or incarcerated before 2030?"* "If China invades Taiwan before 2035, will the US respond with military force?"* "Will there be a tsunami that kills at least 50,000 people before 2030?"* "Will there be a military conflict resulting in at least 50 deaths between the United States and China in 2024?"* "Will an AI system be reported to have successfully blackmailed someone for >\$1000 by EOY 2028?"* "Will Vladimir Putin declare Martial Law in at least 3/4 of Russia before 2025?"Again, when in doubt, do NOT flag the question; mark it as "ok".Your response should take the following structure:Insert thinking:{{ insert your concise thoughts here }}Classification:{{ insert "flag" or "ok"}}
Figure 10: Question Validation Prompt
@@ Appendix G Statistical Details
Appendix H Leaderboards: Top 50
We show the best 50 performers on several leaderboards. The first is the leaderboard for the human question set of 200 standard (non-combination) questions:Table 18. Table 19 shows the leaderboard for the full LLM question set of 1,000 questions. Finally, in Table 20, we present the human leaderboard with combination questions included where humans provided forecasts on both components of the combination question. We derive human forecasts for these combination questions by treating each component of the question as independent.
Table 18: Leaderboard: Human question set (Top 50)